OpenStudy (anonymous):

Compare and Contrast: Two equations are listed below. Solve each equation and compare the solutions. Choose the statement that is true about both solutions. Equation 1 Equation 2 |5x + 6| = 41 |2x + 13| = 28 Equation 1 has more solutions than equation 2. Equation 1 and Equation 2 have the same number of solutions. Equation 2 has more solutions than Equation 1. The number of solutions cannot be determined.

OpenStudy (anonymous):

I am not sure. @phi @dougiee what subject is this?

OpenStudy (phi):

|5x + 6| = 41 if 5x+6 is positive, then this is the same as the equation 5x+6=41 5x = 41-6 (add -6 to both sides) 5x= 35 (now divide both sides by 5) x= 7 if 5x +6 is negative, and we multiply by -1, it will be positive, and we can drop the absolute value operation: -(5x+6) = 41 -5x - 6 = 41 -5x = 47 x = -47/5 we can test both solutions for x in the original equation and see that they are correct

OpenStudy (phi):

try to solve |2x + 13| = 28

OpenStudy (anonymous):

algebra 1 im credits behind,im like far behind being far behind smh:((